OpenStudy (anonymous):

What is the value of z for the equation fraction 1 over 4z = −fraction 7 over 8 + fraction 1 over 8z? −3 −7 3 7

OpenStudy (anonymous):

@lolacole12

OpenStudy (anonymous):

@AlexandervonHumboldt2

zepdrix (zepdrix):

\[\large\rm \frac{1}{4z}=-\frac{7}{8}+\frac{1}{8z}\]This is our problem?

zepdrix (zepdrix):

To make things easier on us, we would like to get rid of the denominators. We can do that by multiplying both sides by the `Least Common Multiple` of our denominators.

zepdrix (zepdrix):

Do you remember how to find LCM? :)

OpenStudy (anonymous):

b

zepdrix (zepdrix):

Hmm the answer should be z=-1/7 are you sure you posted the question correctly? These answer choices don't match up.